1) A 25.0g sample of acetic acid is in a container at STP. What volume is the container?
2) How many molecules of acetic acid are in the container in Problem 1?

Grade:9

1 Answers

Vikas TU
7 years ago
First part => moles = 25/60 = > 0.416
Volume at STP = > 0.416 x 22.4 => 9.33 litres.
 
Second part => 
No. of molecules = > 0.416 x 10^23 x 6.023 = > 2.509 x 10^23